Calculate the sum between the terms
t16 and t53 of an arithmetic sequence with
first term 543 and
common difference –12.
Answer #1

First term (a) = 543


Common difference (d) = –12

nth term of an arithmetic sequence can be calculated using the following formula:

tn = a + (n - 1)d

t16 = 543 + (16 - 1) x - 12

= 543 - 180

= 363

t53 = 543 + (53 - 1) x - 12

= 543 - 624

= - 81

Sum of t16 and t53 = 363 - 81

= 282